# Finite Dynamical Laminations

### v1

@@ Abstract
We developed several combinatorial notions about these laminations some with clear implications for parameter space. We introduce a simplified class of laminations called finite dynamical laminations (FDL). In order to count FDL, we introduce sibling portraits, of which we provide a comprehensive counting theorem. We introduce the pullback tree, and show that all FDL have a child in it. This gives a characterization of which periodic polygons appear in invariant laminations. The base of the pullback tree is a set of laminations, and we show that those laminations have very nice properties. We define the generational FDL graph, and it provides a summary of the combinatorial information we provide about polynomial parameter space.
This work was supported by a UAB Honors College Presidential Honors Fellowship. The author’s research advisor was Dr. John C. Mayer.
@@ 1 Introduction
Laminations have been used to develop an exceedingly, powerful and detailed model of the Mandelbrot set. However, it is much harder in higher degree. In the case of cubics, there is uncertainty about the shape of the set of polynomials associated with the empty lamination. Indeed, the sheer dimension of the set of polynomials makes it difficult to understand the relationships between the polynomials of invariant laminations. Therefore, we introduce a set of simple laminations with a straightforward connection to parameter space, and we aim for a less detailed description of parameter space.
@@ 2 Pullback Trees
With this stipulation, we establish the meaning of the terms parent and child. The natural question is whether all FDL have children. Answering this is a prerequisite for contemplating any polynomials.
@@
3 Dissecting an FDL
Definition 3.1.
A round gap is a component of the complement of a lamination that has arcs of the circle in its boundary. The empty lamination contains one round gap.
Definition 3.2.
A polygon is the boundary of a convex hull of a finite equivalence class of the circle.
The complement of an FDL with the unit circle contains only polygons and round gaps. Lone leaves are polygons, and the term gap always includes them.
Definition 3.3.
A gap, $G$, is of degree $i$ iff $\sigma_{d}|_{Bd(G)}$ is a positively oriented covering map of degree $i$. All lone leaves in an FDL are assigned degree 1. When the image is a leaf, say the leaf from $a$ to $b$, then having degree $i$ means both that there are $2i$ vertices of the polygons and that the polygon alternates between the preimages of $a$ and $b$.
Definition 3.4.
Given a round gap $R$ of degree $i$ and a polygon $P$ contained in $\sigma_{d}(R)$, a sibling portrait is a set of disjoint polygons inside $R$ that map to $P$ and which together have $i$ sides mapping to each side of $P$.
In order to get from one lamination to its child, and be able to see that we never fail to find a child, we consider populating each round gap with the preimages of each polygon in its image. We do this one round gap at a time, and in each round gap we pull back one polygon at a time. But we decide the preimages of any one polygon in any one round gap all at once.
Lemma 3.5.
Given two FDL, $\mathcal{L}$ and $\mathcal{L}^{\prime}$ such that $\mathcal{L}$ is the parent of $\mathcal{L}^{\prime}$, there exists a chain of laminations: $\mathcal{L}_{0}=\mathcal{L},\mathcal{L}_{1},\mathcal{L}_{2},...\mathcal{L}_{n}=L^{\prime}$ such that for each $i$ from 1 to n, the polygons of $\mathcal{L}_{i}\backslash\mathcal{L}_{i-1}$ are a sibling portrait of a polygon of $\mathcal{L}$ in a round gap of $L_{i-1}$.
Proof.
The definition of sibling portrait allows anything allowed in an FDL or invariant lamination. To find such a sequence, first choose an ordering of the polygons in $\sigma_{d}(L^{\prime}\backslash L)$. Form a subsequence $L_{i_{k}}$ of the desired sequence by adding polygons grouped by their image. Next form the actual sequence by adding polygons of $L_{i_{k+1}}$ to $L_{i_{k}}$ in groups by the which round gap of $L_{i_{k}}$ they are in. ∎
Lemma 3.6.
All sibling portraits touch all the preimages of the original polygon’s vertices, and all polygons of a sibling portrait have an equal number of vertices mapping to any particular vertex of the original polygon. Also, no sibling portrait contains any leaf forbidden in an FDL.
Proof.
Since polygons are disjoint even at their endpoints, each preimage vertex is part of at most two sides of one polygon. If the original polygon has an order of vertices like $abc$ then each $b$ must be used the maximal number of times in preimages of $\overline{ab}$ and $\overline{bc}$. If we have a 2-gon, then we have an order $ab$ and in that case, each vertex must be used in two sides of a polygon which map to the two sides of the original polygon. Consider a walk around the sides of a polygon. Since we cannot afford to visit vertices in an order other than that of the original polygon, we only do just that. This excludes not only leaves whose image are interior leaves, but also critical leaves. Moreover, we can only return to our starting place if we are at the end of a cycle. ∎
Lemma 3.7.
All polygons of a sibling portrait have a degree.
Proof.
Assume without loss of generality that the vertices of the original polygon, $P$, are $abcd$. Consider adding one polygon, $P^{\prime}$, of the sibling portrait. Since the $P$ has leaves $ab$, $bc$, $cd$, and $da$; $P^{\prime}$ must also. Therefore, as we walk around the leaves of $P^{\prime}$, we find vertices in an order of either $abcd$ or $dcba$. Since a polygon is a convex hull, the order of its vertices that we get by walking its leaves is the same as the order that they appear on the circle. Suppose that we start drawing $P^{\prime}$ one leaf at a time starting with $a$, and draw the leaves in CCW, counterclockwise, order. In all of $R$, we have the preimages of the vertices in order like $abcdabcd\dots$. Suppose that we then connect from $a$ to $d$ by choosing the reversed order of vertices. Then the arc under that leaf would contain $b$ and $c$ the start and end, so that the order is always $b\dots cdabcdab\dots c$. Since by the previous lemma, a sibling portrait must connect those lone $b$ and $c$ to an $a$ and $b$ without crossing $P^{\prime}$, we cannot finish this sibling portrait. ∎
Lemma 3.8.
If $R$ is a round gap with degree $i$, and $P$ is a polygon contained in $\sigma_{d}(R)$ then the round gaps resulting from inserting a sibling portrait into $R$ have a degree. Moreover, for each piece of the boundary of $\sigma_{d}(R)$ as partitioned by $P$, $i$ pieces of the boundary of $R$ remain unbroken by the preimages of $P$ and map to the same image as before.
Proof.
Consider an original polygon with some order, say, $abcd$. In all of $R$, we have the preimages of the vertices in order like $abcdabcd\dots$. By the previous lemma, up to starting point and number of cycles, all polygons of the sibling portrait will have an order like $abcd$. While we construct a sibling portrait, we add one polygon at a time. Each polygon divides what remains of the circle into arcs if the arc is of degree one, then of course, we have everything we could want. If we have a critical arc, then there are some preimages of the original vertices littered in the arc.
The parent polygon of our critical arc must have, up to starting point, an order $abcd$. Without loss of generality, let the start of the arc be $a$. Thus its end is $b$. Moreover, the first preimage of a vertex appearing in the arc is also $b$. We can moreover note that the sub-arc at the beginning and ending of the critical arc are siblings as well as being consecutive in the circular order of the arc. We then add a polygon starting with $b$ since that is the first preimage vertex on the interior of the arc, and thus the last point of the polygon is $a$ by the polygon’s circular order. In this way, as we remove from the arc by adding polygons, what remains is a number (at least 2) of sibling arcs of the circle. These arcs and leaves of the parent and child polygons form the boundary of a round gap. From the parent polygon, we take the leaf $ab$ and from every child polygon, we also take the leaf $ab$. Therefore, the round gap that is formed maps conjugate to $\sigma_{i}$. ∎
Note that the next two proofs use the same term “directly continuable” as a contraction of two different inductive hypotheses.
Lemma 3.9.
If an FDL contains any preperiodic leaves, it must have at least one child in the pullback tree, and all of its gaps have a degree.
Proof.
A lamination, $L$, is directly continuable if all gaps have a degree and for each gap $G$, there is a collection of gaps $H_{j}$ whose image, $\sigma_{d}(H_{j})$, is equal to $G$ or contains $G$, and such that the total degree of all $H_{j}$ is $d$. Being directly continuable is a sufficient condition for an FDL having children.
By Lemma 3.8, inserting a sibling portrait into a directly continuable lamination produces a directly continuable lamination, and by Lemma 3.5 this means that all children of a directly continuable FDL are also directly continuable.
The empty lamination is directly continuable. Thus, we can infer that the children of the root of the pullback tree are directly continuable because there is no issue with a sibling portrait including its image among its polygons. Therefore, by induction, all FDL containing any preperiodic leaves are directly continuable. Thus, it will have children.
It is easy enough to see that all gaps of an FDL that is not the root have a degree. This means that all polygons of any FDL have a degree. Recall that having a degree is defined as mapping as a positively oriented covering map.
Theorem 3.10.
All FDL have at least one child in the pullback tree.
Proof.
By Lemma 3.9, all that remains to show is that any FDL containing only periodic leaves has at least one child. For the purpose of this proof, all laminations are FDL containing only periodic leaves or the empty lamination. A looser notion than mapping as a covering map is needed to understand these unfortunate laminations. The term covers is defined slightly differently for round gaps and polygons:
  1. 1.
    A round gap, $R$, covers a gap, $G$, $n$ times iff every point in $G\cap\mathbb{S}$ has $n$ preimages in the boundary of $R$ and those preimages are in circular order such that if they are visited in CCW order, then their images would walk around $G\cap\mathbb{S}$ CCW $n$ times.
  2. 2.
    A polygon covers its image once. (Note that periodic polygons are never critical.) A polygon covers all other gaps 0 times.
Unlike in the previous proof, a lamination is directly continuable iff for each gap $G$ in the lamination, there is a collection of gaps {$H_{j}$} such that $H_{j}$ covers $G$ $n_{j}$ times, and $\sum_{j}n_{j}=d$. As in the previous proof, being directly continuable is a sufficient condition for an FDL having children, and the empty lamination is directly continuable.
[…]
The image of $P_{1}$ is $P_{2}$. Let the vertices of $P_{1}$ be labeled $\alpha_{1}$$\alpha_{m}$ in positional order. Let the vertices of $P_{2}$ be labeled $\beta_{i}=\sigma_{d}(\alpha_{i})$, which by the properties of the periodic polygon, is also a CCW positional order.
It is obvious that the pieces of $R$ have enough preimages of $G^{\prime}\cap\mathbb{S}$ in total. To see that those preimages are separated in a good way and are still in the right order we only need to consider 2 cases. The illustration of each case is shown below it.
  1. 1.
    Assume for the moment that $G^{\prime}\neq P_{2}$. Since the bottom picture is still part of an FDL, $G^{\prime}\cap P_{2}$ is at most a leaf. Thus, there is some $i$ such that $G^{\prime}\cap\mathbb{S}\subset[\beta_{i},\beta_{i+1\pmod{m}}]$. Consider the walk around $R\cap\mathbb{S}$. The corresponding walk of $\sigma_{d}(R)$ touches all of $G\cap\mathbb{S}$ in order $n$ times. Starting at $\alpha_{i}$, by the time we get to $\alpha_{i+1\pmod{m}}$, in the image we crossed all of $[\beta_{i},\beta_{i+1\pmod{m}}]$ at least once.
    Consider the siblings of $\alpha_{i}$ and $\alpha_{i+1\pmod{m}}$, which we label $a_{1}$$a_{n}$ and $b_{1}$$b_{n}$ where both are labeled in CCW order and $b_{1}$ is the next $b$ after $a_{1}$. According to the assumption that $R$ covers $G$ $n$ times, $a_{j}$ and $b_{j}$ alternate and the walk from $a_{j}$ to $b_{j}$ corresponds to a satisfactory walk of the arc from $\beta_{i}$ to $\beta_{i+1\pmod{m}}$. Since the $\beta$’s are in circular order, no other $\beta$’s can appear between $\beta_{i}$ and $\beta_{i+1\pmod{m}}$, so no other $\alpha$ angle can appear in between $a_{j}$ and $b_{j}$. Since each walk of $G^{\prime}$ found in the walk of $R$ remains wholly in some piece of $R$, $G^{\prime}$ is covered.
  2. 2.
    The case that $P_{2}$ is equal to $G^{\prime}$ is somewhat similar. We need one fewer arc from $R$ to cover $G^{\prime}$. To ensure that those $n-1$ intervals are unbroken, we label each $\alpha_{i}=a_{i1}$ and each sibling of $\alpha_{i}$ that is less than $\alpha_{i+1\pmod{m}}$ is labeled $a_{ij}$ in order. In that same arc, the siblings of $\alpha_{i+1\pmod{m}}$ are labeled as the alternating $b_{ij}$’s. In this case, since $R$ covers $G$, if any preimages of a vertex of $G^{\prime}$ is appears between $\alpha_{i}$ and $\alpha_{i+1\pmod{m}}$, there must be a complete set of them starting with $b_{ij}$ and ending with $a_{ij}$ which is the opposite of before. In this case also, it is clear that the right number of preimages of $G^{\prime}$ are still in circular order and unbroken by the $\alpha$’s.
The following result is not referenced in this text, and so the reader can be assured that no circular reasoning arises. We hope that it piques the readers interest in further sections. Of course, it has been known that the condition below is a necessary condition for a set of periodic polygons to be found in an invariant lamination [1], but until now it was seemingly unknown that this was a sufficient condition.
Corollary 3.11.
Every pairwise disjoint, forward invariant set of orientation preserving, periodic polygons is in some invariant lamination.
Proof.
Such a set of polygons forms an FDL. By the above theorem, we can make a sequence to laminations by walking down the pullback tree. Later, we call this a sequence to the end of the pullback tree. By Lemma 5.6, such a sequence converges, and the limit contains the starting lamination. By Theorem 5.12, the limit is a proper invariant lamination. ∎
4 Counting Sibling Portraits
Theorem 4.1.
$f(i,n)=\frac{{ni\choose i}}{(n-1)i+1}$
Proof.
Theorem 4.2.
$F(i,n)=f(i,n+1)$
Proof.
We form the inverse of $B$. By Lemma 3.7, all the polygons of a sibling portrait map as a positively oriented covering map. Consider a polygon from an element of the range. It is in the form $abcabcabc\dots$ in positional order. In the whole circle there are multiple $d$ points that we could add, but we are forced to add the ones appearing right before the $a$ vertices of this polygon. (Expanding the polygon in that way cannot cause things to cross since it is the $d$ directly before the $a$). Next we need to break this into polygons, and the only way to do this consecutively ending with $d$. Suppose by way of contradiction that there is some $a$ vertex that we are allowed to connect a $d$ vertex to other than the $a$ preceding it, and finish that polygon as $P_{1}$. Consider an arc of the circle partitioned by that polygon and the vertices in that arc. At least one arc will have a sequence of vertices that does not start with $a$, and the set of polygons formed in that arc will not be placed with $P_{1}$ by $B$. Thus, there is exactly one sibling portrait of the domain for each of the range. ∎

### v2

@@ Abstract
We developed several combinatorial notions about laminations some with clear implications for parameter space. We introduce a simplified class of laminations called finite dynamical laminations (FDL). In order to count FDL, we introduce sibling portraits, of which we provide a comprehensive counting theorem. We provide a characterization of which periodic polygons appear in invariant laminations. We introduce the pullback tree. The base of the pullback tree is a set of laminations, and we show that those laminations are proper and invariant and have a restriction on its critical leaves, so all laminations in the base of the pullback tree correspond to a polynomial. We define the generational FDL graph, and it provides a summary of the combinatorial information we provide about polynomial parameter space.
This work was supported by a UAB Honors College Presidential Honors Fellowship. The author’s research advisor was Dr. John C. Mayer. Some revisions were made under the advice of Dr. Lex G. Oversteegen.
@@ 1 Introduction
Laminations have been used to develop an exceedingly powerful and detailed model of the Mandelbrot set. However, it is much harder in higher degree. In the case of cubics, there is uncertainty about the shape of the set of polynomials associated with the empty lamination. Indeed, the sheer dimension of the set of polynomials makes it difficult to understand the relationships between the polynomials of invariant laminations. Therefore, we introduce a set of simple laminations with a straightforward connection to parameter space, and we aim for a less detailed description of parameter space.
This document explores the uses of finite laminations for studying polynomials. Section 2 introduces the terminology for discussing such laminations, and proves which finite laminations are part of an invariant lamination. In Section 3, we count sibling portraits, which we introduced in Section 2. In Section 4, we introduce Finite Dynamical Lamination, or FDL, a definition which helps us extend combinatorial questions and helps us study the structure of parameter space. Section 5 discusses the limits of sequences of FDL, and Section 6 proves which that all FDL are realized by a polynomial. Section 7 finally shows how to assemble graphs that represent parts of parameter space.
2 The Critical Chords Problem
Given a finite set of disjoint, periodic polygons, when does there exist a $\sigma_{d}$-invariant lamination that contains them? From [4], it is known that the question is the same as whether there is a set of $d-1$ disjoint critical leaves for that set of polygons. Our question is answered in [5], but the argument is longer than necessary. Unlike that document, we will find the critical chords directly, and leave it to the reader to consider how to create an invariant lamination using those leaves.
Definition 2.1.
Given a gap, its basis is its intersection with the circle, and its holes are the circle arcs complementary to its basis.
Definition 2.2.
A subset of the circle is critical if it maps non-injectively by $\sigma_{d}$. A set in the disk (a gap or chord) is critical if its basis is critical.
A $\sigma_{d}$-invariant lamination should be thought of as a $d$ to $1$ invariant lamination. The reader may use either the definition of invariant lamination found in [1] or in [4]. Less than invariant, a lamination may be called forward invariant meaning that the lamination contains its $\sigma_{d}$-image. The next theorem is well known. In addition to proposition II.4.5 of [4], Dr. Oversteegen has promised to write up a proof.
Theorem 2.3.
Given $d-1$ critical leaves compatible with a forward invariant lamination, there exists an invariant lamination containing the original that is also compatible with the critical chords.
Definition 2.4.
A round gap is the closure of a component of the complement of a lamination that has arcs of the circle in its boundary. The empty lamination contains one round gap.
Definition 2.5.
A polygon is the convex hull of a finite, non-singular equivalence class of the circle.
A gap is meant to be a component of the complement of the union of a lamination with the unit disk. The gaps of a finite (one with finitely many leaves) lamination are defined as its polygons and round gaps. Lone leaves are polygons, and the term gap always includes them.
Definition 2.6.
A gap, $G_{i}$, is of degree $d_{i}$ if $\sigma_{d}|_{Bd(G)}$ is a positively oriented covering map of degree $d_{i}$. If the image of $G_{i}$ is a point then $d_{i}$ is the size of $G_{i}$’s basis. When the image is the leaf from $a$ to $b$, then having degree $d_{i}$ means that there are $2d_{i}$ vertices of the polygons and that the polygon’s vertices alternate between the preimages of $a$ and $b$.
Since we mostly think about polygons of a lamination, as opposed to its leaves, and those polygons have siblings, we give a name to a set of sibling polygons. Since a round gap with a degree is like the whole disk, this definition can be with reference to a round gap or the whole disk. We define sibling portrait in a way that is not consistent with the existing notion of sibling collection: a sibling portrait of a polygon $P$ is not, in general, composed of $P$’s siblings.
Definition 2.7.
Given a round gap $R_{i}$ of degree $d_{i}$ and a polygon $P$ contained in $\sigma_{d}(R_{i})$, a sibling portrait is a set of disjoint polygons inside $R_{i}$ that map to $P$ and which together have $d_{i}$ sides mapping to each side of $P$.
Since invariant laminations have siblings for every polygon, the question of whether we can create an invariant lamination comes down to whether we can find places for the siblings. That is, each polygon must be in some sibling portrait that could be added to the lamination. We work to show when this is possible. The next few proofs rely on the order of vertices in a polygon. This is the counterclockwise order of the vertices. When indexing the vertices of an n-gon, subscript is implied to be modulo n.
Lemma 2.8.
A sibling portrait touches all the preimages of the original polygon’s vertices, and all polygons of a sibling portrait have an equal number of vertices mapping to any particular vertex of the original polygon.
Proof.
Since polygons are disjoint even at their endpoints, each preimage vertex is, at most, part of two sides of one polygon. If the original polygon has an order of vertices like $x_{1}x_{2}x_{3}\dots x_{n}$ then each $x_{k}$ must be used the maximal number of times in preimages of $\overline{x_{k-1}x_{k}}$ and $\overline{x_{k}x_{k+1}}$. If we have a 2-gon, then we have an order $ab$ and in that case, each vertex must be used in two sides of a polygon which map to the two sides of the original polygon. Consider a walk around the sides of a polygon. We cannot afford to visit vertices in an order other than that of the original polygon. This excludes not only leaves whose image are interior leaves, but also critical leaves. Moreover, we can only return to our starting place if we are at the end of a cycle. ∎
Remark 2.9.
The preimages of the vertices of a polygon are evenly spaced throughout the circle in the same repeating order as the polygon. (To see this, consider pulling back the polygon according to an all critical d-gon; a half open arc under any side of the d-gon maps monotone onto the circle.) For this reason, there is always at least on sibling portrait that can be formed by using all the vertices in one polygon. If a polygon maps with a degree (with positive orientation) then its siblings are in the same repeating order as itself.
This condition was first noted to be necessary in [1]:
Lemma 2.10.
Given a polygon, $P$, that maps as a covering map, it divides its siblings into even groups iff it maps with positive orientation. In this case, there exists a sibling portrait containing $P$ in the disc.
Proof.
Label the vertices of the original polygon, $\sigma_{d}(P)$, as $x_{1}x_{2}x_{3}\dots x_{n}$. Since the $\sigma_{d}(P)$ has a leaf from each $x_{k}$ to each $x_{k+1}$; $P$ must also. Therefore, as we walk around the leaves of $P$, we find vertices in a, possibly repeating, order of either $x_{1}x_{2}x_{3}\dots x_{n}$ or $x_{n}x_{n-1}x_{n-2}\dots x_{1}$. Since a polygon is a convex hull, the order of its vertices that we get by walking its leaves is the same as the order that they appear on the circle.
Suppose that we start drawing $P$ one leaf at a time starting with $x_{1}$, and draw the leaves in CCW, counterclockwise, order. In all of $R_{j}$, we have the preimages of the vertices in order like $x_{1}x_{2}x_{3}\dots x_{n}x_{1}x_{2}x_{3}\dots x_{n}\dots$. Suppose that we then connect from $x_{k}$ to $x_{k-1}$ while moving CCW by making $P$ orientation reversing. Then the open arc under that leaf would contain $x_{k+1}$ and $x_{k-2}$ the start and end respectively. Using the repeating order of vertices in the arc, we see that it contains fewer $x_{k}$ vertices than $x_{k+1}$. Thus, we can not form complete groups of vertices as needed in the previous lemma.
Alternately, if $P$ is orientation preserving, then under the side from $x_{k}$ to $x_{k+1}$, the first letter in the open arc is $x_{k+1}$ and the last letter in the open arc is $x_{k}$. Thus, under that side, there is some whole number of cycles. ∎
The following lemma is the contribution of Dr Oversteegen. The proof does not require that $\overline{xy}$ is non-crossing with $\overline{ab}$, but it does require that they not be siblings.
Lemma 2.11.
Given a leaf $\overline{ab}$ and a leaf $\overline{xy}$ such that the image $\sigma_{d}(\overline{xy})$ is disjoint from $\sigma_{d}(\overline{ab})$, either side of $\overline{xy}$ has the same number of siblings of $a$ as it does of $b$.
Proof.
Consider that the siblings of $a$,$b$,$x$,$y$ are evenly spaced throughout the circle in a fixed order. We label the siblings of these 4 points without distinguishing one sibling from another. If the $x$ or $y$ are equal to some sibling of $a$ or $b$, then their images obviously cannot be disjoint. Assume without loss of generality that $x$ is in some arc bounded by $a$ and then $b$. If $y$ is in an interval bounded by a sibling of first $b$ and then $a$, then the repeating order of the 4 points is $axby\dots$. A half open arc starting and ending with $a$ will map monotone onto the circle. Thus, the image of $\sigma_{d}(\overline{xy})$ crosses $\sigma_{d}(\overline{ab})$. ∎
Lemma 2.12.
Given a disjoint, finite, forward invariant set of orientation preserving polygons, there exists a lamination containing those polygons such that all polygons are part of a sibling portrait in the disk.
Or, more generally, Given a disjoint, finite set of orientation preserving polygons such that for every pair of polygons, $P_{1}$ and $P_{2}$, either $\sigma_{d}(P_{1})=\sigma_{d}(P_{2})$ or $\sigma_{d}(P_{1})$ is disjoint from $\sigma_{d}(P_{2})$; then there exists a lamination containing those polygons such that all polygons are part of a sibling portrait in the disk.
Proof.
As long as there is a polygon, $P$, left that is not part of a sibling portrait, we can add a polygon that is its sibling. Consider a sibling $a^{\prime}$ of a vertex of $P$. $a^{\prime}$ is on the boundary of a round gap, call it $R$. Every sibling of a vertex $P$ is either in an open arc of $R$ or in the closure of one of its holes. These holes fall into one of 2 categories: either they are bounded by a leaf that is part of a polygon sibling to $P$, or it isn’t. In the first case Lemma 2.10 applies to that polygon; thus, on the side of the polygon containing $R$, there is a complete set of vertices. Using Lemma 2.11, we can make the same conclusion about the other type of hole. Each closure of a hole contains an equal number of siblings of each vertex/letter. Since each vertex has the same number of siblings in the whole disk, each vertex has the same number of siblings in the union of the open arcs of $R$.
The vertices in the open arcs of $R$ are in the same positive order as $P$. Supposing the opposite, then some vertices are skipped. But since they are in positive order in the disk, some hole must contain the skipped vertices. In order to skip a vertex, you have to separate it without putting the other vertices with it. Thus, some closure of a hole of $R$ contains an uneven number of vertices in contradiction to the argument above.
Since those vertices are together in one gap and positively oriented, we can form a new polygon that is also positively oriented. ∎
Lemma 2.13.
  1. i)
    If $R_{j}$ is a round gap with of degree $d_{j}$, and $P$ is a polygon contained in $\sigma_{d}(R_{j})$ then the round gaps resulting from inserting a sibling portrait into $R_{j}$ each have a degree.
  2. ii)
    Consider the set of gaps (including polygons) $\{G_{i}\}$ formed when inserting that sibling portrait into $R_{j}$, and their degrees ${d_{i}}$. Then
    $\sum_{i}(d_{i}-1)=d_{j}-1$
  3. iii)
    Moreover, given a lamination that consists of finitely many disjoint polygons such that each polygon is part of a sibling portrait in the whole disk, all gaps have a degree and the equation above holds. In this case $d=d_{j}$.
Proof.
  1. i)
    The new polygons are already guaranteed to have a degree by Lemma 2.10, so consider only the round gaps created by inserting a sibling portrait into $R_{j}$. Consider an original polygon with some order, say, $x_{1}x_{2}x_{3}\dots x_{n}$. In all of $R_{j}$, that same order repeats. By the previous lemma, up to starting point and number of cycles, all polygons of the sibling portrait will have an order like $x_{1}x_{2}x_{3}\dots x_{n}$. While we construct a sibling portrait, we add one polygon at a time. Each polygon divides what remains of the circle into arcs. If the arc is of degree one, then of course, we have everything we could want. If we have a critical arc, then there are some preimages of the original vertices littered in the arc.
    By Lemma 2.10 after some “first” polygon is inserted, a critical arc created must have, up to starting point, have the same repeating order as $R_{j}$. Let the start of the arc be $x_{k}$. Thus, its end is $x_{k+1}$. Moreover, the first preimage of a vertex appearing in the open arc is also $x_{k+1}$. We can also note that the sub-arc at the beginning and ending of the critical arc are siblings as well as being consecutive in the circular order of the arc. We then add a polygon starting with $x_{k+1}$ since that is the first preimage vertex of the open arc, and thus the last point of the polygon is $x_{k}$ by the polygon’s circular order. In this way, as we remove from the arc by adding polygons until there are no critical arcs, what remains is a number (at least 2) of sibling arcs of the circle. These arcs and leaves of the first and subsequent polygons form the boundary of a round gap. From the parent polygon, we take the leaf $\overline{x_{k}x_{k+1}}$ and from every child polygon, we also take the leaf $\overline{x_{k}x_{k+1}}$. Since the arcs in-between the leaves of this gap all have the same end points, they are also siblings. Therefore, the round gap that is formed has some degree equal to the number of polygons in its boundary.
  2. ii)
    Fix some sibling portrait. Once again, choose a “first” polygon and focus on one side. The degree of the round gap immediately under the first polygon is one greater than the number of polygons immediately under that side of the first polygon. The degree minus 1 of a polygon is the number of extra sets of vertices that it touches. Under each polygon below the first polygon, the subsequent round gaps have degrees that are one greater than the number of polygons immediately below them. According to this counting, each polygon adds degree to the round gap above them. The total $\sum_{i}(d_{i}-1)$ of all the gaps under one side of the first polygon is the number of sets of vertices under that side of the polygon. Since the first polygon takes up one set of vertices besides its own excess degree, the total over all sides and its own excess degree is one less because of that polygon.
  3. iii)
    Consider working up to the lamination inductively, but one sibling portrait at a time. The statement is true of the empty lamination. At each step while inserting a sibling portrait, in-between other sibling portraits, we can see it as inserting a sibling portrait into a round gap with say degree $d_{j}$. Since a round gap with a degree is like the disk, the equation holds for the new gaps inside the round gap. Since the new terms of the sum are equal to what they replace, the equation holds.
The formula above implies something comparable to the Riemann-Hurwitz relation found in [6] when it is applied to the polygons and the round gaps of the lamination. That is the criticality trapped in the polygons plus the criticality free in the round gaps adds to $d-1$.
After forming the sibling portraits, we can place $d-1$ critical chords (without forming a polygon out of them) into gaps in at least one way. Any point in the basis of a critical gap could be the starting point of such a concatenation of critical chords. These critical chords separate the circle into $d$ branches of the inverse which by Theorem 2.3 creates an invariant lamination containing the original set of polygons.
Theorem 2.14.
Given a forward invariant set of disjoint polygons, they are in at least one invariant lamination iff all the polygons map forward as positively oriented covering maps.
It is further asked whether the critical chords can be chosen to touch the original polygons. Consider sliding a critical chord to one side or the other until it can’t slide any further without hitting one of the original polygons. Eventually it will stop as long as the lamination is non-empty.
Another question is whether there exists an invariant lamination that contains every periodic orbit in some list and no more periodic polygons. This is obviously not true. For one example, in degree 3, take $L$ to be the rotational triangles of _001 and _112. The leaf from _0 to _1 is forced. In the original lamination, the round gap where the leaf is forced is not critical. For an example in degree 2, take $L$ to be the orbit of the leaf from _000101 to _101000. A rotational triangle is forced in the non-critical part of the critical gap. Perhaps there is some meaningful sense in which nothing is forced in a non-critical part of the original set.
Conjecture 2.15.
Given a disjoint set of positively oriented periodic polygons, there exists an invariant lamination such that all the additional periodic polygons are of lower period than the originals.
Now consider generalizing Theorem 2.14 by (temporarily) expanding the definitions involved. First recall that a lamination is defined to include all points of the circle. A forward invariant lamination could contain a wandering triangle or a critical leaf with a pre-periodic end point or any number of other things. A finite forward invariant set has finitely many leaves (degenerate leaves don’t count).
Theorem 2.16.
Given a finite forward invariant lamination, it is a subset of at least one invariant lamination iff each polygon in the forward invariant set either has a point as its image or maps as a positively oriented covering map.
Proof.
Take the interpretation that a point of the circle counts as a polygon. Of course, it has no sides, so when we talk about “all sides”, we make a vacuously true statement. Without changing the definition of sibling portrait (Definition 2.7), the sibling portrait of a point can be a set of points. It could also include a critical leaf. Thus, a critical leaf is always part of a sibling portrait in a lamination. The definition of degree (Definition 2.6) plays nicely with this rereading of the term polygon. All the lemmas of the paper hold just as well with, at worst, an added case in the proof. Theorem 2.3 requires the stipulation that a critical chord is compatible with a lamination that contains it. This interpretation is traditional under the convention that compatible means the union is a lamination.
Conjecture 2.17.
The word finite in the theorem above is unnecessary.
3 Counting Sibling Portraits
Theorem 3.1.
$f(i,n)=\frac{{ni\choose i}}{(n-1)i+1}$
Proof.
Theorem 3.2.
$F(i,n)=f(i,n+1)$
Proof.
We form the inverse of $B$. By Lemmas 2.8 and 2.10, all the polygons of a sibling portrait map as a positively oriented covering map. Consider a polygon from an element of the range. It is in the form $abcabcabc\dots$ in positional order. In the whole circle there are multiple $d$ points that we could add, but we are forced to add the ones appearing right before the $a$ vertices of this polygon. (Expanding the polygon in that way cannot cause things to cross since it is the $d$ directly before the $a$). Next we need to break this into polygons, and the only way to do this consecutively ending with $d$. Suppose by way of contradiction that there is some $a$ vertex that we are allowed to connect a $d$ vertex to other than the $a$ preceding it, and finish that polygon as $P_{1}$. Consider an arc of the circle partitioned by that polygon and the vertices in that arc. At least one arc will have a sequence of vertices that does not start with $a$, and the set of polygons formed in that arc will not be placed with $P_{1}$ by $B$. Thus, there is exactly one sibling portrait of the domain for each of the range. ∎
4 Finite Dynamic Laminations
@@
With this stipulation, we establish the meaning of the terms parent and child. We see that all FDL have children from Theorem 2.14.